OpenStudy (brydenwright):

Find all solutions to the equation. sin x = sqrt3/2

OpenStudy (vuriffy):

Oh god.

OpenStudy (vuriffy):

I could help, but let me open my good ole notebook.

OpenStudy (vuriffy):

Oh, I already have this problem done.

OpenStudy (vuriffy):

Do you have a calculator by chance?

OpenStudy (brydenwright):

Yes

OpenStudy (vuriffy):

Okay, for sin x = sqrt(3)/2 sin^-1 each side sin^-1(sin x) = sin^-1 (sqrt(3)/2) Tell me what you get.

OpenStudy (brydenwright):

1.047

OpenStudy (vuriffy):

You have your calculator in the wrong format, actually.

OpenStudy (brydenwright):

so 60?

OpenStudy (vuriffy):

Yes; what is the value of radians at 60?

OpenStudy (brydenwright):

pi/4

OpenStudy (vuriffy):

pi/4 is for 45. pi/3 is for 60.

OpenStudy (brydenwright):

Oh, that's right. Sorry about that.

OpenStudy (vuriffy):

You're fine. So, one of your solutions is pi/3.

OpenStudy (vuriffy):

Also, what I learned by myself is: the value of x times the lower number of the = to value is always the amount of solutions. so (1) - x * (2) - sqrt(3)/2 2 solutions.

OpenStudy (vuriffy):

Sin of x, only does one loop on the above axis, so the solutions are both before 1 pi.

OpenStudy (brydenwright):

\[{ \pi/3+n \pi | n=0, +-1, +-2...}\] is one of the options, would that be it then?

OpenStudy (vuriffy):

Next, you shall use the method of the graph to make it easier for you. All you need to do is PI - pi/3 and you get 2pi/3, which is the second solution.

OpenStudy (brydenwright):

Okay. I think I get it now. Thanks!

OpenStudy (vuriffy):

That was considering the fact of the sin x graph as a whole and also that the rate of distance was pi/3. You're welcome.